# Linear Singer-Hopf conjecture

### v1

@@ 3.2. Some properties of $\Phi_{\eta}$
Let $\eta$ be a multi-valued closed one-form on a complex manifold $X$. Is there a natural lifting of $\Phi_{\eta}$ to a functor of constructible complexes $\Phi_{\eta}:D^{b}_{c}(X,\mathbb{C})\to D^{b}_{c}(X,\mathbb{C})$?
@@
Let $X$ be a smooth projective variety and let $\rho:\pi_{1}(X)\to\mathrm{GL}(r,K)$ be a linear representation where $K$ is a field of characteristic $p>0$. If $\rho$ is large, then there exists a closed mutivalued one-form $\eta$ on $X$ such that for any positive-dimensional closed subvariety $Z$ of $X$, the restriction of $\eta$ to $Z|_{\operatorname{reg}}$ is non-trivial.
@@ 4.1. Positivity from the period maps
Let $X$ be a projective manifold and let $\rho:\pi_{1}(X)\to\mathrm{GL}(r,\mathbb{C})$ be a large and semisimple representation. Then there exists a semisimple representation $\sigma:\pi_{1}(X)\to\mathrm{GL}(N,\mathbb{C})$ on $X$ underlying a $\mathbb{C}$-VHS, and a closed multivalued one form $\eta$ on $X$ such that for any irreducible subvariety $Z$ of $X$, one of the following two statements holds.
@@ 5.2. Techniques from the linear Shafarevich conjecture
  • and a multivalued closed holomorphic one form $\eta$ on $X$
@@
Let $s_{{\rm fac},r}:X\to S_{{\rm fac},r}$ be the factorization map defined in Definition 2.11. By the same arguments in the proof of Proposition 4.12, there exists a closed mutivalued one-form $\eta$ on $X$ such that for any closed subvariety $Z$, $\eta|_{Z_{\rm reg}}$ is trivial if and only if $s_{{\rm fac},r}(Z)$ is a point. From now on, we will assume that $s_{{\rm fac},r}(Z)$ is a point.
@@
We apply Proposition 5.7 to construct a semisimple representation $\sigma_{1}:\pi_{1}(X)\to\mathrm{GL}(N_{1},\mathbb{C})$ underlying a $\mathbb{C}$-VHS, a representation $\sigma_{2}:\pi_{1}(X)\to\mathrm{GL}(N_{2},\mathbb{R})$ underlying a $\mathbb{R}$-VMHS of weight $-1,0$, and a multivalued one form $\eta$ on $X$ such that they satisfies the properties therein. We use the vanishing cycle functor $\Phi_{\eta}$ and apply Corollary 3.10 repeatedly so that we reduce the proof to the case where $\eta|_{Z}$ is trivial. Then the properties in Proposition 5.7 are fulfilled. We then apply Proposition 5.1 to conclude that $T_{Z}^{*}X\cdot T_{X}^{*}X\geq 0$. ∎

### v2

@@ 3.2. Some properties of $\Phi_{\eta}$
Let $\eta$ be a multivalued closed one-form on a complex manifold $X$. Is there a natural lifting of $\Phi_{\eta}$ to a functor of constructible complexes $\Phi_{\eta}:D^{b}_{c}(X,\mathbb{C})\to D^{b}_{c}(X,\mathbb{C})$?
@@
Let $X$ be a smooth projective variety and let $\rho:\pi_{1}(X)\to\mathrm{GL}(r,K)$ be a linear representation where $K$ is a field of characteristic $p>0$. If $\rho$ is large, then there exists a closed multivalued one-form $\eta$ on $X$ such that for any positive-dimensional closed subvariety $Z$ of $X$, the restriction of $\eta$ to $Z|_{\operatorname{reg}}$ is non-trivial.
@@ 4.1. Positivity from the period maps
Let $X$ be a projective manifold and let $\rho:\pi_{1}(X)\to\mathrm{GL}(r,\mathbb{C})$ be a large and semisimple representation. Then there exists a semisimple representation $\sigma:\pi_{1}(X)\to\mathrm{GL}(N,\mathbb{C})$ on $X$ underlying a $\mathbb{C}$-VHS, and a closed multivalued one-form $\eta$ on $X$ such that for any irreducible subvariety $Z$ of $X$, one of the following two statements holds.
@@ 5.2. Techniques from the linear Shafarevich conjecture
  • and a multivalued closed holomorphic one-form $\eta$ on $X$
@@
Let $s_{{\rm fac},r}:X\to S_{{\rm fac},r}$ be the factorization map defined in Definition 2.11. By the same arguments in the proof of Proposition 4.12, there exists a closed multivalued one-form $\eta$ on $X$ such that for any closed subvariety $Z$, $\eta|_{Z_{\rm reg}}$ is trivial if and only if $s_{{\rm fac},r}(Z)$ is a point. From now on, we will assume that $s_{{\rm fac},r}(Z)$ is a point.
@@
We apply Proposition 5.7 to construct a semisimple representation $\sigma_{1}:\pi_{1}(X)\to\mathrm{GL}(N_{1},\mathbb{C})$ underlying a $\mathbb{C}$-VHS, a representation $\sigma_{2}:\pi_{1}(X)\to\mathrm{GL}(N_{2},\mathbb{R})$ underlying a $\mathbb{R}$-VMHS of weight $-1,0$, and a multivalued one-form $\eta$ on $X$ such that they satisfies the properties therein. We use the vanishing cycle functor $\Phi_{\eta}$ and apply Corollary 3.10 repeatedly so that we reduce the proof to the case where $\eta|_{Z}$ is trivial. Then the properties in Proposition 5.7 are fulfilled. We then apply Proposition 5.1 to conclude that $T_{Z}^{*}X\cdot T_{X}^{*}X\geq 0$. ∎